Two asteroids are in space separated by a distance of 4000m with no other forces around them. The first asteroid has a mass of 2000kg and the second one has a mass of 1500kg. a) What is the force of gravity between these two objects?
Answer #1

From Newton's Gravitational Law

The gravitational force between two objects is proportional to their mass and inversely proportional to the square of the distance between their centers.

F =G*m1*m1/(r^2)

F =(6.67*10^-11 N.m^2/kg^2*2000 kg*1500 kg)/(4000 m)^2

F =1.250625*10^-11 N

F =1.250*10^-11 N